OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (ivettef365):

first find the slope on the formula y = 4x + 7

OpenStudy (anonymous):

:/ i dunno how i still figuring out algebra

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

A. y = 4x + 10 B. y = one fourthx + 5 halves C. y = -4x - 6 D. y = -one fourth x + 3 halves

OpenStudy (ivettef365):

ok this is how you read the formula, is important that you know y = mx + b where (x,y) are just points in the line m = slope b = y-int (that means where the lines touches the y)

OpenStudy (ivettef365):

so now you can tell me what is the slope on the formula y = 4x + 7

OpenStudy (anonymous):

Since line is parallel to y = 4x + 7. therefore comparing it with y = mx +c we get slope m = 4 since line passes through the point (-2, 2) there fore x1= -2 and y1 = 2 Now by using the slope p[oint form of the line i.e. y - y1 = m(x-x1) we get y - 2 = 4(x+2) y - 2 = 4x+8 y - 4x = 2+8 y - 4x = 10 i.e. 4x -y + 10 = 0